OpenStudy (anonymous):

how to sketch theta=4pi/3

OpenStudy (anonymous):

theta in regards to what?

OpenStudy (anonymous):

\[\theta=\frac{ 4\pi }{ 3 }\]

OpenStudy (anonymous):

its a sign that looks like a zero with a line in the middle of it

OpenStudy (anonymous):

not sure what else ur looking for brit

OpenStudy (anonymous):

i was wondering how to sketch that on a circle

OpenStudy (anonymous):

ur only pointing to 240 degrees

OpenStudy (anonymous):

That exactly what i couldnt figure out, thank you sooooo much! But how would i figure that out based on the equation i posded earlier?

OpenStudy (anonymous):

you need to memorize the unit circle its vital for passing trig, theres a pattern to it so dont think of it as memorizing everything

OpenStudy (anonymous):

okay thanks.

OpenStudy (anonymous):

the unit circle?

OpenStudy (theviper):

Click on equation than type :- LARGE{color{red}{theta=frac{4pi}{3}}}\[ \LARGE{\color{red}{\theta=\frac{4\pi}{3}}}\]

OpenStudy (anonymous):

LARGE{color{red}{theta=frac{4pi}{3}}}

OpenStudy (anonymous):

ahhhhhhhh seee what you mean